The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.

Answer by checkley71(8403)

2(4X-8)+X=29 OR 8X-16+X=29 OR 9X=45 OR X=45/9 OR X=5 (BASE) THUS
4*5-8=12 (LEGS) BASE + 2 LEGS 5+2*12=29 OR 5+24=29 OR 29=29